OpenStudy (anonymous):

I am trying to calculate the following using a graphing calculator, but cannot seem to come up with the correct answer. PLEASE HELP 8! / 4! 2! 1! 1! (the answer is 840, but how do I come up with this answer?

OpenStudy (debbieg):

Well, just for ease of entering into the calculator, I would not bother with the 1!'s, since they =1. I also wouldn't bother with the "!" on the 2! - that's just 2*1 so = 2. So the expression is really: \[\Large \dfrac{ 8! }{4!2!1!1! }=\dfrac{ 8! }{4!2 }\]

OpenStudy (debbieg):

Now are you using the factorial function on the calculator? or just doing it "manually"? If using the !, just make sure you use ( )'s appropriately: \[\Large 8!/(4!2)\] should get the right answer. Or alternatively: \[\Large \dfrac{ 8! }{4!2 }=\dfrac{ 8\cdot7\cdot6\cdot5}{2 }=8\cdot7\cdot3\cdot5\] and just punch it up!

OpenStudy (anonymous):

Thanks Debbie, the second option seems to work best for me. I need a training session on my graphing calculator!! lol thx again